The price of a particular product is £5.00 per unit, and on average 150 units of the product are sold per month. If the price is increased to £8.00 per unit, demand is expected to fall to an average of 80 units per month. The price elasticity of demand for the product is


1
Expert's answer
2021-11-07T17:03:59-0500

price elasticity of demand for the product =% change in quantity demanded/%change in price

P1 = 5, P2 = 8. Q1 = 150units, Q2 = 80 units

"PED =\frac{Q_2-Q_1}{Q_1} ×100 ÷\frac{P_2-P_1}{P_1} ×100"


"\%\Delta quantity =\frac{80-150}{150}×100\% = -46.67"


"\%\Delta price=\frac{8-5}{5}×100\% =60"


"\because" "PED =-46.67\u00f760 =-0.78"